Calculus II, Spring 2016 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 03

1. Section 10.6: Problems 29, 35, 39, 43, 49, 53,

2. Section 10.7: problems 7, 11, 15, 19, 23, 37, 29.

3. Find a power series that converges on (1, 3) and diverges otherwise. Do the same for (1, 3], [1, 3) and [1, 3], respectively.